# Hegel and Modern Topology

### v6

@@ Abstract
In this paper we sketch how some fundamental concepts of modern topology (as well as logic and category theory) can be understood philosophically in the light of Hegel’s Science Logic as well how modern topological concepts can provide concrete illustrations of many of the concepts and deductions that Hegel used. Also these modern concepts can in turn be very powerful hermeneutic tools permitting a more rigorous and thorough grasp of Hegelian concepts. This paper can be seen as a continuation of our paper [14] where we argued that the prototypes of many fundamental notions of modern topology were already found in Aristotle’s Physics. More generally it is hoped that this note makes a case for the possibility of a rigorous enriching interaction and mutual support between philosophy on one hand and modern logic and mathematics on the other. This paper is obviously meant only as a preliminary sketch and to offer some motivation for exploring in a more detailed and thorough way the subjects discussed. (January 10th: added section Essence and Metaphysics and section Dialectic of the Combination of Systems).
@@
Another overview of the Logic of Being
@@
On Being, Quality and Quantity
@@
Measure from Hegel to Thom
@@
Further considerations on Quantity and Measure
The rich connections between Hegel and Kant, Fichte, Schelling, Leibniz and Spinoza have been substantially explored. However there are lesser known and perhaps even more philosophically interesting connections to Sextus Empiricus (as evidenced in Hegel’s Lectures on the History of Philosophy and of course in the Phenomenology itself), Hume and Proclus. The same goes for the various currents of Buddhist philosophy (which also have deep connections to Hume and Pyrrhonism).
In particular there are some very interesting connections to the theory of dialectics and the Proclean view of mathematics in his Commentary of the the first book of the Elements.
@@
Just as a fragment of IPC is isomorphic to CL, so too CPS style terms can be seen as programs with control (because continuators (which are like the state of the stack) are always passed as arguments).
We recommend F. Agestam’s Interpretations of Classical Logic using $\lambda$-calculus.
@@
Dialectic of the combination of systems
Given two axiomatic-deductive systems S1 and S2, each is an island unto itself, each exists in the world of its own rules, the symbols of which are meaningful only relative to these rules and to the system as a whole. Thus S1 and S2 are neither mutually consistent nor non-consistent - and we can combine them into a new system in a free way. To speak of consistency we must first identify symbols of both systems, we must somehow postulate a common language. And we need a notion of contradiction. But all this only makes sense relative to a third system S3 which acts as an arbiter. If S1 and S2 deduce contradictory things regarding a symbol A, how can we formalize this in S3 ?
By introducing a symbol for S1 and S2 inside the combined system S1++S2, qualifying symbols representing ’from the perspective of S1’ and ’from the perspective of S2’. The introduction of subscripts or symbols for indexing by S1 and S2 is a projection of the system inside itself, a reflection-into-self.
But then S1 and S2 must be qualified by the arbiter system S3. Why is the perspective of S1 relating to A what it is ? (And the same for S2). S1 and S2 share context S3 and the quality of S1 depends on that of S2 and vice-versa. S1 and S2 have become parts of a whole, for-themselves so far as they are for-another. Communication is the outcome of contradiction.
@@
Essence becoming actuality in the Science of Logic: beyond different models for a theory, different frames of reference for physics, representations for an algebraic structure, etc. Do we not here have the phenomenological reflection which neither attempts to have empty thought grasp directly its own structure nor is thought loosing itself in the objectified engagement in its action, but rather is the self-reflected awareness of thought in its thinking, a shift of perspective which knows itself in its process? This goes beyond causality, computation and formal logic to inner and infinite spiritual knowledge. Spiritual knowledge is knowledge that what is in consciousness is taken and proceed from its inmost center, a revelation, veiling-unveiling, which must pass to reveal the process: the revelation is re-velation.
@@
The quest for Notion
@@
Some additional considerations on Hegel’ logic, consciousness and mathematics
@@
About higher topos theory
@@
For 1 and 2 a good reference is Part I of Birgit Richter’s From Categories to Homotopy Theory.
For 3 it is the book chapter Homotopy theories and model categories by W. G. Dwyer and J. Spalinski.
But there is more. One must feel very at home with simplicial sets (and their connection to homotopy types) which play a central role in Higher Topos Theory. Finally one must master the basics of Grothendieck toposes.
Instead of ordinary (set-enriched) model categories we work with model categories enriched over simplicial sets sSet (seen itself as a category enriched over simplicial sets and when endowed with its classical (Quillen) model structure its fibrant-cofibrant objects are called $\infty$-groupoids). Instead of presheaves to sets we work with simplicial set enriched functors from the opposite of a simplicial set enriched category to simplicial sets. A key result is Dugger’s theorem, the simplicial presheaf analogue for combinatorial model categories of the characterization of sheaf toposes as left exact full subcategory localizations of presheaves - itself a generalization of the presentation of an object by generators and relations.
So Higher Topos Theory is Topos Theory done (homotopically) over sSet rather than Set.
@@
Of course there are competing definitions of (models of) $\infty$-groupoids besides the sSet-based one (Kan complexes) which itself is only one possible choice for shapes (which include cubical and cellular sets).
All this suggests philosophically that our different concepts and models of what a ’space’ is are special embodiments of a single ’pure’ concept which is yet to be determined. Note that homotopy type theory connects $\infty$-groupoids to (identity) types/ propositions or spaces of proofs/functions/computations.
On certain concepts of Category Theory
@@
Mereology
A sieve on $x$ is thus a set of parts of $x$ plus all the parts of those parts.
Finally each $z$ is the fusion of all its parts: this is the third axiom
An important aspect of contemporary mereology involves holology, the study of different types and degrees of wholeness in systems. The concept of higher groupoid is relevant to holology. In Martin-Löf type theory there is an extrinsic and an intrinsic concept of equality. The intrinsic concept (written normally $\Gamma\vdash t:A=A$) is what we shall mainly consider here. It can be given a topological (and higher-categorical) interpretation which is of some interest to mereology. Here we only give a rough description of some of the intuitions involved. Consider points in the plane. Internal ’equality’ between two (externally unequal) points $a$ and $b$, written $a=b$, is seen as being witnessed by there being a continuous path from $a$ to $b$. In this way reflexivity, symmetry and transitivity have natural interpretations (in fact we obtain a ’groupoid’ or more correctly a 1-groupoid). Note that there can be more than one path between two points. For a space $C$ consisting of two disjoint ’chunks’ $A$ and $B$ there is no path from a point of $A$ to a point of $B$. There are two extreme situations. One is the discrete case in which no two externally distinct points can be connected by a path - this is the classical concept of a set. The other extreme is a connected (or 1-connected) space in which we can find a single point $p$ such that $p$ can be connected to any other point in the space. Thus the discrete space has the minimum of internal unity and the contractible space the maximum of internal unity. In general we can have a space with a certain number of connected components (’chunks’). If we view internal equality as an equivalence relation (and thus the space as a setoid) then the number of components is the number of equivalence classes.
C. Ortiz Hill’s book[15] contains an interesting discussion on equality and identity in Frege and Husserl. Per Martin-Löf and Frege played similar roles as founders of radical new systems of mathematical logic and foundations of mathematics. As we just saw in Martin-Löf type theory we have two notions of equality. One is extrinsic and corresponds to (absolute) sameness (or identity) and satisfies Leibniz’s law. The other is internal, and corresponds to equality, sameness according to some aspect666it would be interesting to compare this to Aristotle’s statement of the law of non-contradiction in the Metaphysics., which we interpret in terms of paths. Leibniz’s law for internal equality becomes problematic and corresponds topologically to ’liftings’ or ’transports’.
Internal equality itself is equivalent to equivalence.
Aufhebung in Type Theory
Kant and sheaf theory
The sequence seems to be : topos >-> ringed topos >-> lined topos >-> smooth topos. Our task is to elucidate the phenomenological and categorical (in the philosophical sense) meaning of the concept of topos. Also to understand why the internal version of the concepts of commutative ring, commutative algebra, linear map (cf. Weil algebras and the Kock-Lawvere axiom) and an infinitesimal version of simplicial objects (used for defining the dg-algebra of differential forms) appear to be of so fundamental and vast a scope as a priori conditions for mathematical physics and other branches of natural science. And how are we to understand Kant’s concept of space in the light of synthetic differential geometry, it being point-free ? Seeing a vector field on $X$ as an infinitesimal deformation of the identity map on $X$ seems very intuitive.
Thus a category of sheaves becomes interesting and intelligible by its relation to other categories of sheaves. So considering at once the 2-category of sheaf-categories, or more generally, toposes, is very natural and imposes itself in the nature of things. The adjunction between toposes and locally presentable categories is discussed in Marta Bunge, Aurelio Carboni, The symmetric topos, Journal of Pure and Applied Algebra 105:233-249, (1995).
So: section >-> sheaf >-> sheaf category >-> 2-category of sheaf categories. Aquinas (in De Ente et Essentia) views genus as a space of possibilities rather than as a minimal matter to which difference is added as a form. Difference determines or ’picks out’ a latent possibility of the genus. This agrees with the concept of sheaf and category: we choose a section of a sheaf or a definite sheaf in the category of sheaves.
- different orders of wholes (higher groupoids)
Miscellany about Hegel and category theory
Let C be a category representing the totality of individual consciousness and their relations. Then for a given individual consciousness A we have i) the set hom(A,A) (individual, internal relations), ii) the set hom(A,B) relations between two individual consciousnesses, and finally iii) the slice categories (C $\downarrow$ A) and (C $\uparrow$ A) whose objects are all possible relationships between A and its environment (all the other objects in the category C) and whose morphisms are relationships between objects ”from the point of view of A”. Also the process could be repeated, taking slice-categories of slice-categories and so forth.
Elementary toposes represent consciousness and the mind while I think that Grothendieck toposes (toposes of generalized sheaves) are a very abstract representation of physics (and perhaps also systems theory and biology). The underlying site (or in particular topological space or groupoid) represents the most general concept of space-time, the sheaves on it are generalizations of the fields in physics.
Open and dense. This is related to tranversality - again nodes. And to stability, expresssion of mediated quantity invariance of quality.
Topological, open variation, quality, perfection, optimization again involves quantity.
Consider an analogy : day is to night as life is to death. Surely in any analogy there is implicit a correspondence, a set of functions which takes the pairs to each other: F: day,night $\rightarrow$ life,death and G: life,death $\rightarrow$ day,night. Consider the statement: death is the night of life. If this is so then surely: night is the death of day and vice-versa. We take oflifeof\,life and ofdayof\,day to be the functors $F$ and $G$ respectively - here we venture that ’life’ and ’day’ designate the whole genus, the set of both elements of the respective pairs. Thus it would make sense to day : day is the life of day and life is the day of life. And we read this as $death\rightarrow F\,night\Leftrightarrow G\,death\rightarrow night$ which (together with the case for day and life) corresponds to the definition of an adjunction (we can assume we are in a groupoid in which ’is’ is an isomorphism). We can also write symbolically:

### v7

@@ Abstract
In this paper we sketch how some fundamental concepts of modern topology (as well as logic and category theory) can be understood philosophically in the light of Hegel’s Science Logic as well how modern topological concepts can provide concrete illustrations of many of the concepts and deductions that Hegel used. Also these modern concepts can in turn be very powerful hermeneutic tools permitting a more rigorous and thorough grasp of Hegelian concepts. This paper can be seen as a continuation of our paper [14] where we argued that the prototypes of many fundamental notions of modern topology were already found in Aristotle’s Physics. More generally it is hoped that this note makes a case for the possibility of a rigorous enriching interaction and mutual support between philosophy on one hand and modern logic and mathematics on the other. This paper is obviously meant only as a preliminary sketch and to offer some motivation for exploring in a more detailed and thorough way the subjects discussed.
Introduction
Why investigate such a topic as Hegel and Modern Topology? First of all, the author of this paper, acquainted with the major works of Hegel and not being in any sense a ’Hegelian’, proposes a certain framework (the discussion of which must wait for a future paper) for assessing the significance of the Hegel in the history of philosophy with special emphasis on Hegel’s relationship to ancient Hellenic, Hellenistic and Indian philosophy. Doubtlessly such a framework provides some justification or motivation (and this is clear in the context of an Aristotelian[14] or Neoplatonist reading of Hegel) for approaching the topic of Hegel’s relationship to mathematics and more particularly the significance of Hegel for an interpretation of modern mathematics as well as in turn the significance of modern mathematics for an exegesis of Hegel. This is what Van Lambalgen[10] - in a Kantian context - aptly called the ’hermeneutical virtuous circle’.
Studying the mutual interpretation between Hegel’s Science of Logic and modern mathematics is not a new proposal. Hegelian concepts already appear in Lautmann’s philosophy of mathematics. The most advanced developments of modern geometry and topology are inseparable from Category Theory. This last provides a kind of abstract unifying conceptual architecture. It is plauisble that MacLane was conscious of the Hegelian undertones of his use of the tern ’subsumes’ in his treatment of Kan Externsions in his book Categories for the Working Mathematicians. But the first explicit treatment of the relationship of Hegel’s Science of Logic and modern topology and category theory was given by William Lawvere (in the context of his philosophical approach to continuum mechanics and synthetic differential geometry) and there is major contemporary work (by the team responsible for the nLab wesite) in this direction which develops Lawevere’s ideas. This work is focused on homotopy type theory, homotopical algebra and certain attempts at extending topos theory and synthetic differential geometry to a higher category theoretic framework, with notable applications to theoretical physics being carried out by Urs Schreiber.
Hegel’s Logic (both in the Greater Science of Logic and the Encyclopedia Logic) does contain (principally in the Logic of Quantity and Measure) an explicit treatment and critique of the foundations of arithmetic, geometry, infinitesimal analysis as well as applied mathematics. These portitions (or rather Hegel’s extensive comments) have as a rule been dismissed as either inadequate or a merely historical interest (pertaining to a time before the establishment of the modern rigorous foundations for the Calculus). However around the beginning of the last decade interesting research has been carried out by Mascarenhas[1] and Dimitrov[41] offering a more sympathetic reappraisal of Hegel’s approach to mathematics in the Science of Logic and even a partial revindication of it in the light of subsequent developments in the foundations of mathematics.
The goal of this paper is to present a series of sketches regarding how the virtuous hermeneutical circle could be carried out between Hegel’s Science of Logic on one hand and modern topology on the other (as well as category theory and some aspects of mathematical physics, logic and theoretical computer science). These sketches will pertain primarily to the Logic of Being and to the Logic of Essence as we can unfortunately only offer a few suggestions on how the Logic of Concept might be approached.
In this paper we will only refer to basic 20th-century mathematics such as taught in standard undergraduate degrees and in particular remain at the level of ordinary (1-)category theory except some passages where we attempt to give an intuitive explanation for the basic ideas behind homotopy type theory and higher groupoids. But in the appendices we give general (philosophically motivated) introduction to the higher category theoretic and homotopy theoretic approaches found in nLab as well as discussing some more advanced aspects of topos theory and sheaf theory interesting in relationship to Hegel and also in their own right.
The presentation of the material here is obviously very sketchy and its primary aim is to inspire and motivate future research in this area.
@@ Introduction
More on the Logic of Being
@@
The transition from Being through Quality to Quantity
@@
Hegel’s Doctrine of Measure and Thom’s Catastrophe Theory
@@
More on Quantity and Measure
The connections between Hegel and Kant, Fichte, Schelling, Leibniz and Spinoza have been substantially explored. However there are lesser known and perhaps even more philosophically interesting connections to Sextus Empiricus (as evidenced in Hegel’s Lectures on the History of Philosophy and of course in the Phenomenology itself), Hume and Proclus. The same goes for the various currents of ancient Indian philosophy (themselves with interesting connections to Pyrrhonism and Hume). A topic to be explored is Hegel’s relationship to Proclus’ theory of dialectics and philosophy of mathematics in the Commentary of the the first book of the Elements.
@@
Just as a fragment of IPC is isomorphic to CL, so too CPS style terms can be seen as programs with control (because continuators (which are like the state of the stack) are always passed as arguments)111for an introduction to this subject we recommend F. Agestam’s Interpretations of Classical Logic using $\lambda$-calculus..
@@
The Dialectic of the Combination of Systems
Given two axiomatic-deductive systems $S_{1}$ and $S_{2}$, each is an island unto itself, each exists in the world of its own rules, the symbols of which are meaningful only relative to these rules and to the system as a whole. Thus $S_{1}$ and $S_{2}$ are neither mutually consistent nor non-consistent - and we can combine them into a new system in a free way. To speak of consistency we must first identify symbols of both systems, we must somehow postulate a common language. And we need a notion of contradiction. But all this only makes sense relative to a third system S3 which acts as an arbiter. If $S_{1}$ and $S_{2}$ deduce contradictory things regarding a symbol $A$, how can we formalize this in $S_{3}$?
By introducing a symbol for $S_{1}$ and $S_{2}$ inside the combined system $S_{1}+S_{2}$, qualifying symbols representing ’from the perspective of $S_{1}$’ and ’from the perspective of $S_{2}$’. The introduction of subscripts or symbols for indexing by $S_{1}$ and $S_{2}$ is a projection of the system inside itself, a reflection-into-self.
But then $S_{1}$ and $S_{2}$ must be qualified by the arbiter system $S_{3}$. Why is the perspective of $S_{1}$ relating to $A$ what it is ? (And the same for $S_{2}$). $S_{1}$ and $S_{2}$ share context $S_{3}$ and the quality of $S_{1}$ depends on that of $S_{2}$ and vice-versa. $S_{1}$ and $S_{2}$ have become parts of a whole, for-themselves so far as they are for-another. Communication is the outcome of contradiction.
@@
Essence becoming actuality in the Science of Logic: beyond different models for a theory, different frames of reference for physics, representations for an algebraic structure, etc. Do we not here have the phenomenological reflection which neither attempts to have empty thought grasp directly its own structure nor is thought loosing itself in the objectified engagement in its action, but rather is the self-reflected awareness of thought in its thinking, a shift of perspective which knows itself in its process? This goes beyond causality, computation and formal logic to inner and infinitel knowledge. Knowledge that what is in consciousness is taken and proceeds from its inmost center, a revelation, veiling-unveiling, which must pass to reveal the process: the revelation is re-velation.
@@
There appear to be very interesting mutually illuminating connections between Schopenhauer’s very abstract, universal, apodeictic and relational formulation of the principle foufold root of sufficient reason (and its further development in The World of Will as Representation) and Hegel’s Doctrine of Essence (specially the treatment of absolute relation, necessity and actuality). This is a topic worth exploring in future research.
@@
Difficulties in Approaching Concept (Notion)
@@
Notion, Consciousness and Mathematics
@@
Mereology and Homotopy Type Theory
A sieve on $x$ is thus a set of parts of $x$ plus all the parts of those parts.
Finally each $z$ is the fusion of all its parts: this is the third axiom
An important aspect of contemporary mereology involves holology, the study of different types and degrees of wholeness in systems (we repeat here some of the discussion in section 12). The concept of higher groupoid is relevant to holology. In Martin-Löf type theory there is an extrinsic and an intrinsic concept of equality. The intrinsic concept (written normally $\Gamma\vdash t:A=A$) is what we shall mainly consider here. It can be given a topological (and higher-categorical) interpretation which is of some interest to mereology. Here we only give a rough description of some of the intuitions involved. Consider points in the plane. Internal ’equality’ between two (externally unequal) points $a$ and $b$, written $a=b$, is seen as being witnessed by there being a continuous path from $a$ to $b$. In this way reflexivity, symmetry and transitivity have natural interpretations (in fact we obtain a ’groupoid’ or more correctly a 1-groupoid). Note that there can be more than one path between two points. For a space $C$ consisting of two disjoint ’chunks’ $A$ and $B$ there is no path from a point of $A$ to a point of $B$. There are two extreme situations. One is the discrete case in which no two externally distinct points can be connected by a path - this is the classical concept of a set. The other extreme is a connected (or 1-connected) space in which we can find a single point $p$ such that $p$ can be connected to any other point in the space. Thus the discrete space has the minimum of internal unity and the contractible space the maximum of internal unity. In general we can have a space with a certain number of connected components (’chunks’). If we view internal equality as an equivalence relation (and thus the space as a setoid) then the number of components is the number of equivalence classes.
C. Ortiz Hill’s book[15] contains an interesting discussion on equality and identity in Frege and Husserl. Per Martin-Löf and Frege played similar roles as founders of radical new systems of mathematical logic and foundations of mathematics. As we just saw in Martin-Löf type theory we have two notions of equality. One is extrinsic and corresponds to (absolute) sameness (or identity) and satisfies Leibniz’s law. The other is internal, and corresponds to equality, sameness according to some aspect555it would be interesting to compare this to Aristotle’s statement of the law of non-contradiction in the Metaphysics., which we interpret in terms of paths. Leibniz’s law for internal equality becomes problematic and corresponds topologically to ’liftings’ or ’transports’.
Internal equality itself is equivalent to equivalence.
Aufhebung in Type Theory
Kant and Sheaf Theory
The sequence seems to be : topos $\rightarrow$ ringed topos $\rightarrow$ lined topos $\rightarrow$ smooth topos. Our task is to elucidate the phenomenological and categorical (in the philosophical sense) meaning of the concept of topos. Also to understand why the internal version of the concepts of commutative ring, commutative algebra, linear map (cf. Weil algebras and the Kock-Lawvere axiom) and an infinitesimal version of simplicial objects (used for defining the dg-algebra of differential forms) appear to be of so fundamental and vast a scope as a priori conditions for mathematical physics and other branches of natural science. And how are we to understand Kant’s concept of space in the light of synthetic differential geometry, it being point-free ? Seeing a vector field on $X$ as an infinitesimal deformation of the identity map on $X$ seems very intuitive.
Thus a category of sheaves becomes interesting and intelligible by its relation to other categories of sheaves. So considering at once the 2-category of sheaf-categories, or more generally, toposes, is very natural and imposes itself in the nature of things [^6] The adjunction between toposes and locally presentable categories is discussed in Marta Bunge, Aurelio Carboni, The symmetric topos, Journal of Pure and Applied Algebra 105:233-249, (1995)..
So: section $\rightarrow$ sheaf $\rightarrow$ sheaf category $\rightarrow$ 2-category of sheaf categories. Aquinas (in De Ente et Essentia) views genus as a space of possibilities rather than as a minimal matter to which difference is added as a form. Difference determines or ’picks out’ a latent possibility of the genus. This agrees with the concept of sheaf and category: we choose a section of a sheaf or a definite sheaf in the category of sheaves.
- different orders of wholes (higher groupoids)
A Recapitulation
Let $C$ be a category representing the totality of individual consciousness and their relations. Then for a given individual consciousness A we have i) the set $hom(A,A)$ (individual, internal relations), ii) the set $hom(A,B)$ relations between two individual consciousnesses, and finally iii) the slice categories ($C\downarrow A$) and ($C\uparrow A$) whose objects are all possible relationships between A and its environment (all the other objects in the category $C$) and whose morphisms are relationships between objects ”from the point of view of AA”. Also the process could be repeated, taking slice-categories of slice-categories and so forth.
Elementary toposes represent consciousness and the mind while Grothendieck toposes (toposes of generalized sheaves) are a very abstract representation of physics (and perhaps also systems theory and biology). The underlying site (or in particular topological space or groupoid) represents the most general concept of space-time, the sheaves on it are generalizations of the fields in physics.
Open and dense. This is related to tranversality - again nodes. And to stability, expresssion of mediated quantity and invariance of quality. We can also investigate the topological aspects of the calculus of variations.
Consider an analogy : day is to night as life is to death. Surely in any analogy there is implicit a correspondence, a set of functions which takes the pairs to each other:
F: day,night $\rightarrow$ life,death
and
G: life,death $\rightarrow$ day,night
Consider the statement: death is the night of life. If this is so then surely: night is the death of day and vice-versa. We take of life and of day to be the functors $F$ and $G$ respectively - here we venture that ’life’ and ’day’ designate the whole genus, the set of both elements of the respective pairs. Thus it would make sense to day : day is the life of day and life is the day of life. And we read this as $death\rightarrow F\,night\Leftrightarrow G\,death\rightarrow night$ which (together with the case for day and life) corresponds to the definition of an adjunction (we can assume we are in a groupoid in which ’is’ is an isomorphism). We can also write symbolically:
Appendix 1: Homotopy Theory and Higher Categories
@@ A Recapitulation
For 1 and 2 a good reference is Part I of Birgit Richter’s From Categories to Homotopy Theory. For 3 it is the book chapter Homotopy theories and model categories by W. G. Dwyer and J. Spalinski.
One must feel very at home with simplicial sets (and their connection to homotopy types). Instead of ordinary (set-enriched) model categories we work with model categories enriched over simplicial sets sSet (seen itself as a category enriched over simplicial sets and when endowed with its classical (Quillen) model structure its fibrant-cofibrant objects are called $\infty$-groupoids). Instead of presheaves to sets we work with simplicial set enriched functors from the opposite of a simplicial set enriched category to simplicial sets. A key result is Dugger’s theorem, the simplicial presheaf analogue for combinatorial model categories of the characterization of sheaf toposes as left exact full subcategory localizations of presheaves - itself a generalization of the presentation of an object by generators and relations. Higher Topos Theory is Topos Theory done (homotopically) over sSet rather than Set.
@@
Of course there are competing definitions of (models of) $\infty$-groupoids besides the sSet-based one (Kan complexes) which itself is only one possible choice for shapes (which include cubical and cellular sets). All this suggests philosophically that our different concepts and models of what a ’space’ is are special embodiments of a single ’pure’ concept which is yet to be determined. Note that homotopy type theory connects $\infty$-groupoids to (identity) types/ propositions or spaces of proofs/functions/computations.
Appendix 2: aspects of Topos Theory and Sheaf Theory